OpenStudy (anonymous):

What is the graph of the equation? 2 x- y = 2

OpenStudy (anonymous):

2x - y = 2 -2x -2x (subtract 2x) -y = -2x + 2 -1 -1 -1 (divide by -1) y = 2x - 2 slope (m) = 2 y intercept = -2

OpenStudy (anonymous):

What you could also do is find the intercepts which is where you plug in zero for x or y. To find x: 2x - (0) = 2 2x = 2 Divde by 2 on both sides. x = 1